A competitive firm is maximizing profits by producing 250 units of output at the current market price of RM1000 per unit. The firm has average fixed cost (AFC) of RM300 and total costs of RM300000 at this output level.

How to calculate TFC, ATC, AVC, MC, TR, and MR.
1
Expert's answer
2017-02-21T08:35:45-0500
TFC = AFC*Q = 300*250 = $75000
ATC = TC/Q = 300000/250 = $1200
AVC = ATC - AFC = 1200 - 300 = $900
MC = (TC2 - TC1)/(Q2 - Q1), so we need additional info to calculate it
TR = P*Q = 1000*250 = $250000
MR = (TR2 - TR1)/(Q2 - Q1), so we also need additional info to calculate it